OpenStudy (anonymous):

Simplify x+1/2x^2+5x+3 Divided by 2/2x+3

OpenStudy (anonymous):

\[\frac{ x+1 }{2x ^{2}+5x+3 }\div \frac{ 2 }{ 2x+3 }\]

OpenStudy (kc_kennylau):

Factorize \(2x^3+5x+3\) first :)

OpenStudy (babyslapmafro):

\[=\frac{ x+1 }{ 2x ^{2}+5x+3 } \times \frac{ 2x+3 }{2 }\]

OpenStudy (kc_kennylau):

Factorize \(2x^2+5x+3\) first :)

OpenStudy (babyslapmafro):

@kc_kennylau if you multiply by the reciprocal first it allows you to identify a factor that will most likely be cancelled

OpenStudy (anonymous):

you both are confusing me

OpenStudy (kc_kennylau):

Just factorize \(2x^2+5x+3\) first

OpenStudy (anonymous):

@kc_kennylau so it would be 2x(x+5x+3)

OpenStudy (kc_kennylau):

nope it would be \((2x+3)(x+1)\)

OpenStudy (anonymous):

@kc_kennylau so then i would write \[\frac{ x+1 }{ (2x+3)(x+1) }\div \frac{ 2 }{ 2x+3 }\]